OpenStudy (anonymous):

the roots of the equation x^3-4x-12=0 are a,b,c Find the equation with roots a+b, b+c, c+a

OpenStudy (anonymous):

One way would be to find a,b and c

hartnn (hartnn):

sum of roots = a+b+c = -0/1 = 0 so u just have to find and equation whose roots are -c,-a,-b.

OpenStudy (anonymous):

IS IT ALWAYS TRUE @hartnn sum of roots of a cubic equation =0

OpenStudy (anonymous):

i guess a,b,c are not real

OpenStudy (anonymous):

Hmm, I've found the sum and products of the roots but I don't know how to find a,b,and c and @hartnn, how do I find an equation with roots -c,-a and -b?

hartnn (hartnn):

if the equation is Ax^3+Bx^2+cx+d=0, then sum of roots = -B/A is always true

OpenStudy (anonymous):

oh. SINCE b=0

OpenStudy (anonymous):

sum of roots=0

OpenStudy (anonymous):

Ok then this may be a nice approach, Since the roots are -a,-b,-c (-x)^3 -4(-x)-12=0 -x^3+4x-12=0

hartnn (hartnn):

lets see, u have roots as, -a,-b-,c its product = -abc = -(-12) =12 sum = (-a-b-c) = -(a+b+c) = 0 so its of the form : x^3+px+12 =0

hartnn (hartnn):

or what @sauravshakya did.

hartnn (hartnn):

u are correct, the final equation then becomes. x^3-4x+12=0

OpenStudy (anonymous):

YEP

OpenStudy (anonymous):

ok, I got the rest of it but why did the roots have to be -b,-c, -a?

hartnn (hartnn):

u know how to find sum and product of roots of the general equation : Ax^3+Bx^2+Cx+D=0 ?

OpenStudy (anonymous):

yup!

hartnn (hartnn):

sum = -B/A product = D/A now apply this to your equation x^3-4x-12=0 and tell me what u get sum as ?

OpenStudy (anonymous):

sum is -b/a which is 4

OpenStudy (anonymous):

opps I mean -0/4

hartnn (hartnn):

its -0/1 = 0

OpenStudy (anonymous):

opps again that's what I meant

hartnn (hartnn):

so a+b+c=0 a+b=-c b+c=-a a+c=-b

OpenStudy (anonymous):

oh I get it now! Thank you!!!

hartnn (hartnn):

welcome :) that was a good question.